A chunk of nickel weighing 18.0 grams and originally at 97.55 °C is dropped into an insulated cup containing 81.0 grams of water at 23.10 °C.

Assuming that all of the heat is transferred to the water, the final temperature of the water is ______ °C.
Answer #1

specific heat capacity of Nickel = 0.440 J/g.oC

m(water) = 81.0 g

T(water) = 23.1 oC

C(water) = 4.184 J/goC

m(nickel) = 18.0 g

T(nickel) = 97.55 oC

C(nickel) = 0.44 J/goC

T = to be calculated

Let the final temperature be T oC

use:

heat lost by nickel = heat gained by water

m(nickel)*C(nickel)*(T(nickel)-T) = m(water)*C(water)*(T-T(water))

18.0*0.44*(97.55-T) = 81.0*4.184*(T-23.1)

7.92*(97.55-T) = 338.904*(T-23.1)

772.596 - 7.92*T = 338.904*T - 7828.6824

T= 24.8001 oC

Answer: 24.80 oC